Extra help needed for this problem.
A manufacturer employs 24 part-tome workers. This is 30% of their employees. What is their total employee count?


I really need help with this problem can someone explain it to me?

Answer by stanbon(75887)

A manufacturer employs 24 part-tome workers. This is 30% of their employees. What is their total employee count?
------------
Let the total employee count be "x".
EQUATION:
30% of x = 24
0.30x = 24
x = 24/0.3
x = 240/3
x = 80 (the total number of employees is 80)
============
Cheers,
Stan H.